Hello,
How do I solve this by the system by addition. I'm not sure how to go about it.
This is what I get,but my friend says it's wrong. I'm loss fro this point on.
5x – 3y = 13
4x – 3y = 11
+_______________
9x - 6y= 11

Answer by adamchapman(301)

5x – 3y = 13....................1
4x – 3y = 11....................2
rearrange (1):
-5x + 3y = -13..................3
Add (2) and (3):
4x – 3y + (-5x + 3y) = 11 -13 ..................4
-x=-2
x=2
substitute x=2 into (1):
10 - 3y = 13
-3y = 3
y=-1
So the solution is
x=2,
y=-1
I hope this helps,
Adam
P.S. please visit my website, it may be helpful to you. The address is www.geocities.com/quibowibbler